OpenStudy (anonymous):

NEED HELP GEOMETRY! Two similar rectangles have lengths 18 in and 12 in. What is the ratio of their widths? What is the ratio of their perimeters?

OpenStudy (anonymous):

The ratio of their widths will be the same as the ratio of their lengths: L1/L2 = 18/12 = 3/2 The ratio of their perimeters will also be: 3/2 because (2L1 + 2W1)/(2L2 + 2W2) = (L1 + W1)/(L2 + W2)

OpenStudy (anonymous):

would the area be the same ratio?

OpenStudy (anonymous):

@tcarroll010

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@tcarroll010 hi! can you pleasee help me when you are done here?

OpenStudy (anonymous):

sure @erin512

OpenStudy (anonymous):

how would i find it out?

OpenStudy (anonymous):

The ratio of the areas is 9/4 which is (3/2)^2

OpenStudy (anonymous):

thanks so much :)

OpenStudy (anonymous):

why would you square it?

OpenStudy (anonymous):

First area is LW, the second area is [(3/2)L] x [(3/2)W] = (9/4) of the first area. Yiou square it because both the length and width are 3/2 bigger.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

You're welcome and thx for the recognition!